1.) The cost of an asset is $ 1,080, 000​, and its residual value is $ 290,000. Estimated useful life of the asset is five years. Calculate depreciation for the first year using the double-declining-balance method of depreciation.​ (Do not round any intermediate​ calculations, and round your final answer to the nearest​ dollar.

2.)On January​ 1, 2018, Tyson Manufacturing Corporation purchased a machine for $40,600,000. ​Tyson's management expects to use the machine for 28,000 hours over the next six years. The estimated residual value of the machine at the end of the sixth year is $43,000. The machine was used for 3,700 hours in 2018 and 5,100 hours in 2019. What is the depreciation expense for 2018 if the corporation uses the units−of−production method of​ depreciation? (Round any intermediate calculations to two decimal​ places, and your final answer to the nearest​ dollar.)

3.) Orbit​ Services, Inc. pays $760,000 to acquire 30​% ​(200,000 shares) of the voting stock of State​ Investments, Inc. on January​ 5, 2019. State​ Investments, Inc. declares and pays a cash dividend of $2.40 per share on June​ 14, 2019. Which of the following is the correct journal entry for the transaction on June​ 14, 2019?
Answer #1

1) First year Depreciation = 1080000*40% = $432000

2) Depreciation expense = (40600000-43000)/28000 = 1448.46*3700 = $5359302

3) Journal entry

Date account and explanation Debit Credit
June 14 Cash (200000*2.4) 480000
Investment in Stock 480000